Write an equation of a line that passes through the points (-2,1) and it is parallel to line 4X+2Y=-1
Found 2 solutions by mananth, Cromlix:
Answer by mananth(16946)

4 x + 2 y = -1
Find the slope of this line
make y the subject
2 y = -4 x -1
Divide by 2
y = -2 x - 1/2
Compare this equation with y=mx+b
slope m = -2
The slope of a line parallel to the above line will be the same
The slope of the required line will be -2
m= -2 ,point ( -2 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = 4 + b
b= -3
m= -2
Plug value of the slope and b in y = mx +b
The required equation is y = -2 x -3

m.ananth@hotmail.ca

Answer by Cromlix(4381)

Sort 4x + 2y = -1
into y = mx + c form
2y = -4x - 1
y = -2x - 1/2
Lines that are parallel to
one another have the same
slope.
y = -2x - 1/2
has a slope of -2
Using equation:
y - b = m(x - a)with coords (-2,1)
a = -2, b = 1 and m = -2
y - 1 = -2(x -(-2))
y - 1 = -2(x + 2)
y = -2x -4 + 1
y = -2x - 3.
Hope this helps.
:-)